OpenStudy (anonymous):

What exponential function is the best fit for the data in the table? x f(x) 3 10 4 16 5 28 f(x) = 3(2)^(x − 2) + 4 f(x) = 3(2)^(x − 2) − 4 f(x) = 1/3(2)^(x − 2) + 4 f(x) = 1/3(2)^(x− 2) − 4

OpenStudy (anonymous):

@sammixboo

sammixboo (sammixboo):

Let's start with \(\tt \color{red}{f(x)=3(2)^{(x − 2)} + 4}\) We can plug in 3 for\(\tt \color{red}{x}\), and 10 for \(\tt \color{red}{f(x)}\), and see if the equation it true. \(\tt \color{red}{10=3(2)^{(3 − 2)} + 4}\) \(\leftarrow\) Is this true?

sammixboo (sammixboo):

If you are unsure if it is true, we can work it out and see.

OpenStudy (anonymous):

10= 6^(1)+4 10= 6 + 4 10 = 10 YES

sammixboo (sammixboo):

:) So your answer is A

OpenStudy (anonymous):

Thank You Sammi you are really smart!

sammixboo (sammixboo):

Make sure you do the exponent first when solving a problem. It was ok to do the multiplying first, but next time make sure to solve the numbers with the exponents

OpenStudy (anonymous):

Okay will do!